On January 1, 2010, Z Corp (one of only two companies in the world working on New Years Day) purchased a large piece of equipment from Y Corp (the other one) at a cost of $1 million. Z also paid $25,000 to have the thing delivered, $40,000 in labor to have the machine installed and another $35,000 in labor and energy cost to test the machine before it was ready to be used for production. Finally on July 1, 2010 the machine began to be used in production. The machine is expected to last for 10 years and the company does not expect to be able to sell the machine at the end of the 10 year life. How much depreciation expense will be recognized in 2010 and 2011?
Answer #1

Using Straight line method of depreciation

Cost of the equipment = Purchase price+ Delivery+ Installation+ Testing cost

=1000000+ 25000+ 40000+ 35000

=1100000

Annual depreciation = (Cost of equipment- Salvage)/ Useful life

= (1100000-0)/ 10

= 110,000

In 2010: Since machine is ready after half year,

Depreciation claimed will be 110000/2 = 55000

In 2011

Depreciation claimed= 110,000